FOURIER TRANSFORM INFRARED EMISSION SPECTRA OF MgH AND MgH_{2}

\underline{A. SHAYESTEH}, D. R. T. APPADOO, I. GORDON, and P. F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, ON, N2L 3G1, Canada.

We have recorded the vibration-rotation spectra of gaseous MgH and MgH_{2} in emission using a furnace-discharge source. The molecules were generated at 650^{o}C and 333 mA discharge current with magnesium and a mixture of argon and hydrogen gases. The recorded spectra contained several emission bands, as well as the absorption of atmospheric H_{2}O. The highest signal-to-noise ratio for MgH lines was about 200. Three vibrational bands, v=1-0 to v=3-2, for ^{24}MgH and two vibrational bands for ^{25}MgH and ^{26}MgH were observed in the ^{2}Σ^{+} ground electronic state. The analysis of the infrared data combined with our previous data on the B′-X electronic transition will lead to an improved potential energy curve for the ground state using a direct-potential-fit approach. In addition to MgH, we found the antisymmetric stretching mode (ν_{3}) of ^{24}MgH_{2} and three hot bands involving ν_{2} and ν_{3} in our spectrum. The bands were rotationally analyzed and the spectroscopic constants were determined. The MgH_{2} molecule has a linear structure with an R_{0} Mg-H bond length of 1.703327(3) Å.